which element makes up 20.16% of mass NH4CIO


1
Expert's answer
2021-02-01T03:47:00-0500

M=14+4+35.5+16=69.6 g/mol

w=M(E)/M; M(E) = 69,6*0.2016= 14 g/mol = N (nitrogen)